Please help me with all 33... meanthing (A),(B),(C),(D), and (E).
33. (5 pts) What volume in mL of 0 3. (s pts) what volume in mL of 0.500 M solution of HCIO (aą) is needed to neutralize 100 mL of 0.200 M NaOHag) solution? (Start by writing and balancing the reaction.) 0.200 M NaOH(aq) solution? (Start by writing and balancing the (a) 10.0 mL (e) none of these (b) 20.0 mL (c) 40.0 mL (d) 80.0 mL
Answer #1

number of mole = (molarity)*(volume)
since, n- factor of HClO4 and NaOH is 1
number of mole of HClO4 = number of mole of NaOH
0.5*V = 0.2*100
V = (200/5) ml
V = 40 ml

Answer: option C